# 1st Chenyang Xu    2nd Siming Li    3rd Hao Wang*

### v1

@@ I Introduction
Cardiovascular diseases (CVDs) remain the leading cause of mortality worldwide, accounting for approximately 32% of global deaths annually . Early and accurate diagnosis through cardiac auscultation is critical for timely intervention, yet the availability of diverse pathological heart sound recordings remains severely limited, particularly for rare conditions and pediatric populations [oliveira2021circor]. This data scarcity fundamentally constrains the development of robust AI-driven diagnostic systems, such as those for complex multi-task learning like murmur grading and risk analysis [xu2023cardiac], and limits opportunities for medical education and clinical training.
@@
Recent breakthroughs in generative AI, particularly diffusion models [ho2020denoising, rombach2022high] and their application to biosignal synthesis [wang2024medddm], have opened new avenues for addressing medical data scarcity[wassyng2015can]. However, existing approaches suffer from limited controllability, a lack of clinical interpretability, and insufficient validation of medical accuracy. Moreover, simple label-conditioned generation ignores the interplay between different cardiac pathologies and their acoustic manifestations.
@@ II-A Generative Models for Medical and Audio Synthesis
Synthetic data generation is a critical approach to address data scarcity in medicine [mccharty2003]. The field has progressed from classical signal processing to deep generative models, including Variational Autoencoders (VAEs) [teli2025variational] and, more recently, diffusion models. These have been applied to diverse tasks, from ECG synthesis to 3D medical image and shape generation [wang2024medddm].
In parallel, diffusion models have set a new state-of-the-art in audio synthesis [ho2020denoising, dhariwal2021diffusion]. Early models like DiffWave [kong2021diffwave] and WaveGrad [chen2021wavegrad] operated on raw waveforms, while Latent Diffusion Models (LDMs) [rombach2022high] improved efficiency by working in a compressed latent space. This latent-space approach proved highly effective for text-to-audio generation, as demonstrated by systems like AudioLDM [liu2023audioldm], MegaTTS [jiang2025megatts], and LatentSpeech [lou2024latentspeech]. These models typically leverage joint text-audio embeddings, often trained via contrastive learning [radford2021learning], to achieve fine-grained control via natural language. Our work builds upon these foundations, applying a specialized LDM to the clinical domain of PCG synthesis.
@@ II-B Hierarchical Latent Diffusion Models
While sharing the emphasis on multi-level latent priors with Zhang et al. [zhang2024nested], our work distinguishes itself by defining a domain-specific, physiologically-grounded hierarchy rather than a general-purpose one. Similarly, we adapt the principle of disentangled manipulation from Lu et al. [lu2023hierarchical], who focused on images, to the unique temporal and pathological attributes of biosignals. Our model also builds on the paradigm of hierarchical LDMs for structured synthesis, demonstrated by Kim et al. [kim2023neuralfield] for 3D scenes, but we tailor this concept for 1D medical signals by incorporating domain-specific constraints via our Medical Attention module. We also acknowledge other architectural parallels in recent works like LION [zeng2023lion], Hyper-Transforming LDMs [rombach2023hypertransforming], and HiCo [wang2024hico], which collectively underscore the power of hierarchical designs—a trend our work contributes to within the specialized medical domain.
@@ III-A Hierarchical Clinical Metadata Encoding
This structured text $\mathcal{T}_{\text{medical}}$ is encoded by a medically fine-tuned BERT model [alsentzer2019bioclinicalbert]. To capture patient-specific relational data, we process a patient-specific subgraph $\mathcal{G}_{\text{patient}}$ using a 2-layer GraphSAGE network ($g_{\text{GNN}}$). The final conditioning embedding $\mathbf{c}$ is a learned fusion of these two modalities:
@@ IV-A Experimental Setup
Our primary training corpus is the PhysioNet CirCor dataset [oliveira2021circor, reyna2021circor], a highly imbalanced collection of 5,282 recordings from 1,568 pediatric patients. We employed a strict patient-level 70% 15% 15% split for robust evaluation. To assess generalization, we utilized an external dataset of 500 recordings, which was expert-annotated and included diverse pathologies absent in CirCor. Finally, a panel of five pediatric cardiologists with over 10 years of experience was assembled to evaluate the generated signals for clinical validity, physiological plausibility, and educational utility.

### v2

@@ I Introduction
Cardiovascular diseases (CVDs) remain the leading cause of mortality worldwide, accounting for approximately 32% of global deaths annually . Early and accurate diagnosis through cardiac auscultation is critical for timely intervention, yet the availability of diverse pathological heart sound recordings remains severely limited, particularly for rare conditions and pediatric populations [12]. This data scarcity fundamentally constrains the development of robust AI-driven diagnostic systems, such as those for complex multi-task learning like murmur grading and risk analysis [22], and limits opportunities for medical education and clinical training.
@@
Recent breakthroughs in generative AI, particularly diffusion models [4, 15] and their application to biosignal synthesis [19], have opened new avenues for addressing medical data scarcity[21]. However, existing approaches suffer from limited controllability, a lack of clinical interpretability, and insufficient validation of medical accuracy. Moreover, simple label-conditioned generation ignores the interplay between different cardiac pathologies and their acoustic manifestations.
@@ II-A Generative Models for Medical and Audio Synthesis
Synthetic data generation is a critical approach to address data scarcity in medicine [11]. The field has progressed from classical signal processing to deep generative models, including Variational Autoencoders (VAEs) [17] and, more recently, diffusion models. These have been applied to diverse tasks, from ECG synthesis to 3D medical image and shape generation [19].
In parallel, diffusion models have set a new state-of-the-art in audio synthesis [4, 3]. Early models like DiffWave [7] and WaveGrad [2] operated on raw waveforms, while Latent Diffusion Models (LDMs) [15] improved efficiency by working in a compressed latent space. This latent-space approach proved highly effective for text-to-audio generation, as demonstrated by systems like AudioLDM [8], MegaTTS [5], and LatentSpeech [9]. These models typically leverage joint text-audio embeddings, often trained via contrastive learning [13], to achieve fine-grained control via natural language. Our work builds upon these foundations, applying a specialized LDM to the clinical domain of PCG synthesis.
@@ II-B Hierarchical Latent Diffusion Models
While sharing the emphasis on multi-level latent priors with Zhang et al. [23], our work distinguishes itself by defining a domain-specific, physiologically-grounded hierarchy rather than a general-purpose one. Similarly, we adapt the principle of disentangled manipulation from Lu et al. [10], who focused on images, to the unique temporal and pathological attributes of biosignals. Our model also builds on the paradigm of hierarchical LDMs for structured synthesis, demonstrated by Kim et al. [6] for 3D scenes, but we tailor this concept for 1D medical signals by incorporating domain-specific constraints via our Medical Attention module. We also acknowledge other architectural parallels in recent works like LION [18], Hyper-Transforming LDMs [16], and HiCo [20], which collectively underscore the power of hierarchical designs—a trend our work contributes to within the specialized medical domain.
@@ III-A Hierarchical Clinical Metadata Encoding
This structured text $\mathcal{T}_{\text{medical}}$ is encoded by a medically fine-tuned BERT model [1]. To capture patient-specific relational data, we process a patient-specific subgraph $\mathcal{G}_{\text{patient}}$ using a 2-layer GraphSAGE network ($g_{\text{GNN}}$). The final conditioning embedding $\mathbf{c}$ is a learned fusion of these two modalities:
@@ IV-A Experimental Setup
Our primary training corpus is the PhysioNet CirCor dataset [12, 14], a highly imbalanced collection of 5,282 recordings from 1,568 pediatric patients. We employed a strict patient-level 70% 15% 15% split for robust evaluation. To assess generalization, we utilized an external dataset of 500 recordings, which was expert-annotated and included diverse pathologies absent in CirCor. Finally, a panel of five pediatric cardiologists with over 10 years of experience was assembled to evaluate the generated signals for clinical validity, physiological plausibility, and educational utility.